# Distributionally Generative Augmentation for Fair Facial Attribute Classification

### 11 Mar 2024 v1

@@ 2 Related work
Bias Mitigation without Group Information. The labels of spurious attributes may be unavailable in practice. To this end, recently studies have explored how to mitigate bias without additional annotation [21, 2, 81, 6, 40, 41, 3, 71, 78, 57]. Most of these methods predict the bias information as the proxies for the spurious attributes by some heuristics (usually the prediction errors given by a biased classifier) [34, 49, 8, 39]. In this paper, we develop a novel approach to identify the potential spurious attributes via generative techniques. Compared to the previous works, our method perform fairness-aware distributionally generative augmentation, contributing to the quality of the representations. Moreover, our method enhances interpretability by visualizing the detected spurious attributes in image space.
Generative Modeling. Generative models have achieved great success in recent years [42, 24, 11]. InterFaceGAN has shown that different semantics are encoded in the latent space of GANs, and the latent codes are usually well-disentangled after linear transformations [61]. Based on this finding, corresponding techniques for image attribute manipulation have been proposed [65]. Some works have proposed to evaluate fairness by generating counterfactual samples [10, 27, 9]. Recently, generation-based methods have demonstrated significant strides in bias mitigation by constructing a balanced and unbiased dataset [55, 53, 36, 75, 26]. However, these methods need the prior of additional annotations. In this paper, we extend the generation-based approach to cases without additional annotations, further improving fairness and enhancing interpretability.
Representation Learning. Learning representations is important for reliable performance in visual recognition. Recent years, contrastive learning has been remarkably successful in learning effective representations [66, 69, 74, 4, 23, 17, 50, 5]. However, traditional representation learning methods ignore potential fairness issues. To this end, as a pre-processing method, fair representation learning has achieved great success [73, 47, 7, 77, 67, 60, 51, 14]. For example, FSCL proposes to learn fair representations by closing the distance of samples with the same target attribute labels but different sensitive attribute labels [52]. However, most of existing fair representation learning methods rely on labels of spurious attributes, while our method avoids this limitation by the proposed bias detection method.
@@ 4.4 Bias Detection Results on Facial Datasets
Implementations. We perform grid search for the optimal combination coefficients $c^{*}_{1}$, $c^{*}_{2}$ such that the combined direction $c^{*}_{1}\boldsymbol{n}_{1}-c^{*}_{2}\boldsymbol{n}_{2}$ only contains semantics of spurious attributes. The criterion is to make the target attribute match most consistent before and after editing, that is, to make the changes of the target attributes as little as possible. In order to improve efficiency, we randomly sample only 100 images from the training dataset. For each search, we edit them positively and negatively respectively by the combined direction to get 200 edited images. Then we predict their target labels by using CLIP as reference model with prompts “A face with/without smile”. For ease of search, we rewrite the combined direction as $c_{1}(\boldsymbol{n}_{1}-c_{2}/c_{1}\boldsymbol{n}_{2})$. So we only need to search for $c_{2}/c_{1}$ (from 0.5 to 1.5 with unit 0.1). The results are shown in Table 1. We set $c_{1}/c_{2}=0.9$, which makes the target attribute change the least after editing.

### 25 Mar 2024 v2

@@ 2 Related work
Bias Mitigation without Group Information. Some studies have explored how to mitigate bias without additional annotation [22, 2, 87, 6, 43, 44, 3, 76, 84, 60]. Most of these methods predict the bias information as the proxies for the spurious attributes by some heuristics (usually the prediction errors given by a biased classifier) [37, 52, 8, 42]. Recently, researchers have tried to improve robustness based on pre-trained models (e.g., CLIP [57]) without additional annotations [17, 28, 73, 35, 78, 65]. Note that these efforts rely on pre-trained models, and our work can be done without using pre-trained models. For example, we can use a reference model trained by JTT [42] instead of CLIP.
Generative Modeling for Fairness. Generative models have achieved great success in recent years [45, 25, 11, 64, 69]. Some works have proposed to evaluate fairness by generating counterfactual samples [10, 29, 9]. Recently, generation-based methods have demonstrated significant strides in bias mitigation by constructing a balanced and unbiased dataset [58, 56, 39, 81, 27]. However, these methods need the prior of additional annotations. In this paper, we extend the generation-based approach to cases without additional annotations.
Fair Representation Learning. Learning representations is important for reliable performance in visual recognition. Recent years, contrastive learning has been remarkably successful in learning effective representations [70, 74, 80, 4, 24, 18, 53, 5]. However, traditional representation learning methods ignore potential fairness issues. To this end, as a pre-processing method, fair representation learning has achieved great success [79, 50, 7, 83, 71, 63, 54, 14]. For example, FSCL proposes to learn fair representations by closing the distance of samples with the same target attribute labels but different sensitive attribute labels [55]. However, most of existing fair representation learning methods rely on labels of spurious attributes, while our method avoids this limitation by the proposed bias detection method.
@@ 4.3 Bias Detection Results on Facial Datasets
Implementations. We perform grid search for the optimal combination coefficients $c^{*}_{1}$, $c^{*}_{2}$ such that the combined direction $c^{*}_{1}\boldsymbol{n}_{1}-c^{*}_{2}\boldsymbol{n}_{2}$ only contains semantics of spurious attributes. The criterion is to make the target attribute match most consistent before and after editing, that is, to make the changes of the target attributes as little as possible. In order to improve efficiency, we randomly sample only 100 images from the training dataset. For each search, we edit them positively and negatively respectively by the combined direction to get 200 edited images. Then we predict their target labels by using CLIP as reference model with prompts “A face with/without smile". For ease of search, we rewrite the combined direction as $c_{1}(\boldsymbol{n}_{1}-c_{2}/c_{1}\boldsymbol{n}_{2})$. So we only need to search for $c_{2}/c_{1}$ (from 0.5 to 1.5 with unit 0.1). The results are shown in Table 1. We set $c_{1}/c_{2}=0.9$, which makes the target attribute change the least after editing.
@@ 5 Conclusions
Acknowledgment This work was supported by the National Natural Science Foundation of China (62337001, 62376243, 62037001, U20A20387), the Fundamental Research Funds for the Central Universities (No. 226-2022-00051), the StarryNight Science Fund of Zhejiang University Shanghai Institute for Advanced Study (SN-ZJU-SIAS-0010), Project by Shanghai AI Laboratory (P22KS00111) and National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-RP-2021-022). Long Chen is supported by HKUST Special Support for Young Faculty (F0927) and HKUST Sports Science and Technology Research Grant (SSTRG24EG04).